Find the length of the curve described by the polar equation. e = COS 2

please show detailed solution, thank you
Answer #1

NOTE: cuore: -C1) x= f(0) The length of the curve (1) from 8 = d to 0= ß is [fo]: [f@y]? do B I= dra tenta de LSoluno Güven curve: go a los A=0 8-91 0=311/2 dr do - 1/2 sino Now, - = 4 Sin g= costo 021/2 = (26056) = (1+(050) Since, red